OpenStudy (anonymous):

what is -2(p-6)=-14 p=?

OpenStudy (anonymous):

Hey!!! I know you! You helped me sooo many times before! Your awesome!

OpenStudy (texaschic101):

-2(p - 6) = -14 --- divide both sides by -2 p - 6 = -14/-2 p - 6 = 7 p = 7 + 6 p = 13 check.. (-2(13 - 6) = -14 -2(7) = -14 -14 = -14 (correct) so p = 13 any questions ?

OpenStudy (anonymous):

nope! hahah!! Thanks!

OpenStudy (texaschic101):

lol...happy to help :)